# Refining CNN-based Heatmap Regression with Gradient-based Corner Points for Electrode Localization

### v1

@@ Abstract
keywords:
Heatmap regression , Keypoint detection , Electrode localization
journal: arxiv
@@ 1 Introduction
However, due to the multi-levels structure of the cells and the significant overlap between the electrode sheets, X-ray images often suffer from low brightness, high noise, and variations in pole area sizes and positions [2]. These challenges make accurate electrode detection particularly difficult. Existing automated detection algorithms fail to fully leverage the unique features of X-ray images, leading to a high rate of overkill.
@@
  • 4.
    A joint optimization framework combining CNN-based heatmap regression [5] and gradient-based corner point detection is proposed, validated by various evaluation metrics. Additionally, PCS (Percentage of Correct Samples) is introduced as a novel metric to calculate the proportion of samples with maximum normalized error below a threshold.
@@ 2 Methodology
Our framework for electrode localization and optimization is illustrated in the Fig. 1 (a). Given high-resolution X-ray images of the lithium-ion battery, we initially employ a corner point detection model to extract a set of corner points. These corner points are subsequently input into an ROI (Region of Interest) estimation module to identify a smaller area for further focus. Within this ROI, pole positions are detected using a CNN-based heatmap regression model. Finally, the pole positions predicted by the CNN model are refined and corrected by leveraging the corner points previously identified through gradient-based corner detection.
@@ 2.1 Corner Points and ROI Detection
After obtaining the center point, we accumulate the row-wise grayscale values of the X-ray image to form a new ”pattern,” as shown in Fig. 1 (b). This pattern enables us to easily identify the upper and lower bounds of the region of interest (ROI), which correspond to the locations where the row-wise accumulated gradient experiences a significant change. We set a threshold $\tau$ to detect these gradient changes, and define two pointers: the upper pointer moves from 0 to $y_{0}$ with a step size $\Delta b$, while the lower pointer moves from the image height $H$ to $y_{0}$ with a step size $\Delta a$. This process can be formalized as:
@@
where $n$ and $m$ are the step indices, and $\Delta b$ and $\Delta a$ are the step sizes for the upper and lower pointers, respectively. Once we have the upper bound $b$ and the lower bound $a$, we introduce a new variable $\lambda$ to control the width of the ROI region, which is determined by the horizontal coordinate of the cluster center and the height of the ROI region. We therefore achieve the left bound $c$ and right bound $d$ as follows:
@@
where $\lambda$ is a scaling factor that adjusts the width of the region of interest (ROI) based on the cluster center’s horizontal position and the height of the ROI.
@@ 2.2 Electrode Regression Network
Once we have the ROI region, we can use any keypoint detection method to estimate the poles in the image. Due to the significant advantages of HRNet in high-resolution feature extraction and precise keypoint localization, it has been widely adopted for high-accuracy keypoint detection tasks. Additionally, HRNet utilizes heatmap regression to locate keypoints, which enhances the model’s accuracy in keypoint detection. Therefore, we choose HRNet as our electrode regression network.
@@ 2.3 Post-Correction with Corner Points
Before evaluating the confidence of $p_{i},q_{i}$, we first need to establish a ”reference standard” which will be used to assess the affinity of $p_{i}$ and $q_{i}$ to this standard. This affinity represents the confidence information we aim to obtain. Our local feature extractor can be implemented in various ways, such as using BRIEF (Binary Robust Independent Elementary Features) [6] descriptors or other deep learning-based descriptors [7]. Here, we adopt a more efficient approach: for a given point $x_{i},y_{i}$, we define a small square region centered around it, and compute the normalized brightness distribution histogram of this region as its local pattern.
Due to the presence of both positive and negative terminals in lithium batteries, with slight differences in their local patterns, we use two colors in the figure: red for the positive terminal. To address this, we design a standard feature extractor that summarizes the local patterns of both the positive and negative terminals in an input X-ray image, calculating their average as the reference feature $r_{i}$. Similarly, we use a patch feature extractor to extract the local features $u_{i},v_{i}$ for $p_{i}$ and $q_{i}$ individually. Thus the corresponding confidence
@@ 3.1 Experimental setting
We first apply OFAST to detect $N$ corner points. Based on this, we determine the Region of Interest (ROI) and input it into HRNet. Finally, we use our confidence-based coordinate optimization module. We provide baseline results based on HRNet and compare the results with different choices of $N$, corresponding to different numbers of corner points extracted by OFAST.
@@ 3.2 Evaluation Metric
Normalized Mean Error (NME) is used to evaluate the accuracy of predicted keypoints compared to the ground truth. It calculates the mean of the normalized errors across all samples and keypoints, with a lower NME indicating better performance
@@ 4 Conclusion
Acknowledge
Thanks to Z.P.J. and Dr. H.D. for the discussions on this project from Fengchao Energy.

### v3

@@ 1 Introduction
However, due to the multi-levels structure of the cells and the significant overlap between the electrode sheets, X-ray images often suffer from low brightness, high noise, and variations in pole area sizes and positions [2]. These challenges make accurate electrode detection particularly difficult. Existing automated detection algorithms fail to fully leverage the unique features of X-ray images, leading to a high rate of overkill. To this end, we propose a novel joint optimization framework that incorporates a gradient-based corner point detection method to refine CNN-based heatmap regression for electrode localization.
@@
  • Various evaluation metrics validate the effectiveness of our joint optimization framework, which combines CNN-based heatmap regression [5] and gradient-based corner point detection. Additionally, PCS (Percentage of Correct Samples) is introduced as a novel metric to calculate the proportion of samples with maximum normalized error below a threshold.
2 Related Work
Electrode Localization in X-ray Images.
Precise electrode localization is essential for quality assurance in lithium-ion battery manufacturing, particularly in safety-critical applications such as electric vehicles, where misaligned electrodes can pose significant hazards [2, 6]. Traditional methods, including Canny edge detection [7] and Hough transforms [8], assume clear geometric boundaries and fail under the low contrast and heavy overlap characteristic of real-world industrial X-ray images. Recent learning-based approaches have focused on coarse-grained tasks such as defect classification[9] or component segmentation [10], but largely overlook the fine-grained localization required for accurate pole coordinate estimation. Our work bridges this gap by introducing a novel approach for high-precision pole coordinate detection.
Heatmap Regression with Geometric Priors.
Heatmap regression is widely adopted for keypoint detection across domains, from human pose estimation [11] to medical image analysis [12], with HRNet [4] achieving state-of-the-art performance via consistent high-resolution representations. However, two challenges arise in the context of X-ray electrode localization: (1) weak boundary gradients result in diffuse or ambiguous heatmaps, as also noted in facial landmark detection [13]; and (2) conventional CNNs lack awareness of global geometric structure. To overcome these limitations, we propose a hybrid framework that combines learnable heatmap prediction with deterministic corner proposals generated by gradient-based detectors, thereby incorporating explicit geometric priors into the localization process.
Gradient-Driven Corner Detection.
Classical corner detectors such as Harris, FAST [14], and OFAST [3] remain valuable in industrial imaging due to their robustness under noise—an advantage that has been underexploited in recent deep learning pipelines. Unlike segmentation-driven methods [10], we employ OFAST not as a standalone detector, but as a geometric regularizer that guides and refines CNN-based localization. This strategy parallels recent work on hybrid vision systems for document layout analysis [15], but is adapted here to address electrode-specific challenges such as multi-layer interference and occlusion.
@@ 3 Methodology
Our framework for electrode localization and optimization is illustrated in the Fig. 1 (a). Given high-resolution X-ray images of the lithium-ion battery, we initially employ a corner point detection model to extract a set of corner points. These corner points are subsequently input into an ROI estimation module to identify a smaller area for further focus. Within this ROI, pole positions are detected using a CNN-based heatmap regression model. Finally, the pole positions predicted by the CNN model are refined and corrected by leveraging the corner points previously identified through gradient-based corner detection.
@@ 3.1 Corner Points and ROI Detection
After obtaining the center point, we accumulate the row-wise grayscale values of the X-ray image to form a new "pattern," as shown in Fig. 1 (b). This pattern enables us to easily identify the upper and lower bounds of the ROI, which correspond to the locations where the row-wise accumulated gradient experiences a significant change. We set a threshold $\tau$ to detect these gradient changes, and define two pointers: the upper pointer moves from 0 to $y_{0}$ with a step size $\Delta b$, while the lower pointer moves from the image height $H$ to $y_{0}$ with a step size $\Delta a$. This process can be formalized as follows:
@@
where $n$ and $m$ are the step indices, and $\Delta b$ and $\Delta a$ are the step sizes for the upper and lower pointers, respectively. Once we have the upper bound $b$ and the lower bound $a$, we introduce a new variable $\lambda$ to control the width of the ROI, which is determined by the horizontal coordinate of the cluster center and the height of the ROI. We therefore achieve the left bound $c$ and right bound $d$ as follows:
@@
where $\lambda$ is a scaling factor that adjusts the width of the ROI based on the cluster center’s horizontal position and the height of the ROI.
@@ 3.2 Electrode Regression Network
Once we have the ROI, we can use any keypoint detection method to estimate the poles in the image. Due to the significant advantages of HRNet in high-resolution feature extraction and precise keypoint localization, it has been widely adopted for high-accuracy keypoint detection tasks. Additionally, HRNet utilizes heatmap regression to locate keypoints, which enhances the model’s accuracy in keypoint detection. Therefore, we choose HRNet as our electrode regression network.
@@ 3.3 Post-Correction with Corner Points
Before evaluating the confidence of $p_{i},q_{i}$, we first need to establish a "reference standard" which will be used to assess the affinity of $p_{i}$ and $q_{i}$ to this standard. This affinity represents the confidence information we aim to obtain. Our local feature extractor can be implemented in various ways, such as using BRIEF (Binary Robust Independent Elementary Features) [16] descriptors or other deep learning-based descriptors [17]. Here, we adopt a more efficient approach: for a given point $x_{i},y_{i}$, we define a small square region centered around it, and compute the normalized brightness distribution histogram of this region as its local pattern.
Due to the presence of both positive and negative terminals in lithium batteries, with slight differences in their local patterns, we use two colors in the figure: red for the positive terminal. To address this, we design a standard feature extractor that summarizes the local patterns of both the positive and negative terminals in an input X-ray image, calculating their average as the reference feature $r_{i}$. Similarly, we use a patch feature extractor to extract the local features $u_{i},v_{i}$ for $p_{i}$ and $q_{i}$ individually. Thus the corresponding confidence $\alpha_{i},\beta_{i}$ can be achieved as follows:
@@ 4.1 Experimental setting
We first apply OFAST to detect $N$ corner points. Based on this, we determine the ROI and input it into HRNet. Finally, we use our confidence-based coordinate optimization module. We provide baseline results based on HRNet and compare the results with different choices of $N$, corresponding to different numbers of corner points extracted by OFAST.
@@ 4.2 Evaluation Metric
Normalized Mean Error (NME) is used to evaluate the accuracy of predicted keypoints compared to the ground truth. It calculates the mean of the normalized errors across all samples and keypoints, with a lower NME indicating better performance.